Grade 11General Physics

the position of a body is given by x=At+Bt^3 ,where A and B are constants,x is position and t is time.find acc. as a function of time and velocity n acc. at t=5 secs

3:09 pmplss help me to solve it out

vaidehi verma
9 Years agoGrade 11

1 Answer

Sagar jha
9 Years ago
X=at+bt^3 then v=a+3bt^2..... (1) put t=5 in equation 1 uh get v=a+75b and a=6bt then put t=5 uh get a=30b